The sum of 2 numbers is 138. If three times the smaller number is subtracted from the larger number, the result is 22. Find the two numbers. Present your answer as an ordered pair.
Help I know I need two equations but do not know where to begin. What do I do to solve this problem?

Answer by Paul(988)

x+y=138
y=138-x
x-3(138-x)=22
x+3x-414=22
4x=436
x=109
y=138-(109)
y=29
Hence the numbers are 22 and 109